OpenStudy (smurfy14):

how do you do 2^(3x-4)=5 ?? *its a logrithim*

OpenStudy (anonymous):

use: \[\ln (2^{(3x-4)})= \ln5\]

OpenStudy (anonymous):

Take the logarithm of both sides, and the exponent (3x-4) becomes a coefficient because of logarithm rules.\[(3x-4) \log(2) = \log(5)\] \[\rightarrow 3x-4 = \log(5)/\log(2) \rightarrow x=\log(5)/(3\log(2))+4/3\]

OpenStudy (smurfy14):

after the second arrow..how did you get x = log(5)/(3log(2))+4/3? did you divide by three or something?

OpenStudy (anonymous):

Yeah, I skipped a step, sorry. xD I just divided each term by 3.

OpenStudy (smurfy14):

so the answer should be x=4.774 right?